OpenStudy (anonymous):

what is the smart solution of If 3! = 3x2x1, calculate the value from 1 x 1! + 2 x 2! + 3 x 3! + ... + 9 x 9!

OpenStudy (raden):

i love the number "9", just look the last sequence the answer will be = (9+1)! - 1

OpenStudy (shubhamsrg):

1 x 1! + 2 x 2! + 3 x 3! + ... + 9 x 9! = (2-1) 1! + (3-1)2! + (4-1) 3! ........ (10-1) 9! = (2! + 3! .. 10!) - ( 1! +2! ... 9!) = as @RadEn said..

OpenStudy (anonymous):

thank's for help :)